Find a fraction between 1/4 and 1/3 whose numerator and denominator add up to 47. Thanks for any help.
Answer by Alan3354(69443)

Find a fraction between 1/4 and 1/3 whose numerator and denominator add up to 47.
------------
1/4 <= n/d <= 1/3
3 <= 12n/d <= 4
n + d = 47
------------
3 <= 12n/(47-n) <= 4
141 - 3n <= 12n <= 188 - 4n
141 + n <= 12n <= 188
n <= 12n <= 47
11n <= 47 - n
--> n = 11
d = 36
--> 11/36 & 10/37 are the 2 integer solutions.